### v1

@@ 4.1 Experiment Settings
Evaluation Metrics. We evaluate retriever model quality using 4 standard IR (Information Retrieval) metrics. These are (i) NDCG@K/N@K (Normalized Discounted Cumulative Gain): evaluates the quality of a ranked list by rewarding relevant items being placed at the top of the list and penalizing them being placed towards the bottom; (ii) P@K (Precision): The fraction of retrieved items that are relevant to the search; (iii) R@K (Recall): The fraction of relevant items that were retrieved, and (iv) MRR (Mean Reciprocal Rank): Inverse rank of the first relevant item in the retrieved list. For each measure, @K means the number of items that are retrieved. Following ToolRet, we set K to 10.
Baselines. In §2, we mention various related approaches for tool-retriever training. However, Xu et al. (2024) does not have publicly available code. As such, we use, (i) Shi et al. (2025) (ToolRet (TR)), who train their models with basic query-tool similarity; (ii) COLT Qu et al. (2024), a complex two-step framework that first tunes a model (Contriever Izacard et al. (2022)) using standard InfoNCE with query-tool similarity, followed by further optimisation with a graph neural network (GNN) using various decompositions of the training dataset; (iii) use HT directly for retrieval similar to Kachuee et al. (2025) (K in Table 1).
HT Generation Model. We use GPT-4.1 OpenAI (2025) to generate hypothetical tools for both training and test data queries. Additionally, we show that ToolDreamer can also use open-source models for HT generation (§4.3.1). Thus, it is not restricted by the choice of LLM used.
4.2 Tool Retrieval Performance Comparison
To evaluate the tool retrieval performance of ToolDreamer, we choose a dense Qwen3-8B and a sparse BM25 model (App. A). Table 1 shows the results of ToolDreamer along with related baselines.
@@ 4.2 Tool Retrieval Performance Comparison
Finally, we see how strong our framework is against COLT, another related baseline. Firstly, the results from phase-1 training (general query-tool alignment) are far superior to phase-2 (GNN-optimisation). This makes sense as they train their model using the standard query-tool InfoNCE objective, which works relatively well. However, the increased complexity of grouping tools, aligning them with queries, and using a GNN for further optimisation plummets performance. In this regard, ToolDreamer is a much simpler and scalable approach capable of handling a variety of questions.
4.3 Ablation Studies
We conduct various ablation studies to investigate the importance of each step of ToolDreamer: (i) the impact of HT quality - also, the effect of switching the HT-generator to an open-source model (§4.3.1) (ii) using a weak discriminator for tool alignment along with the alignment algorithm itself (§4.3.2), and (iii) whether using an LLM for ranked list fusion offers additional benefits or not (§4.3.3).
@@ 4.3.1 Hypothetical Tool Quality
Different Hypothetical Tool Generator. In our experiments, we use GPT-4.1 to create our HT’s. However, this might be a bottleneck in cases where API-calling might be an issue, either due to cost or privacy concerns Yao et al. (2024). Thus, we wanted to see how switching out GPT-4.1 for an open-source model impacted performance. To this end, we use Qwen3-32B (with the same prompt (Fig. 8) as GPT-4.1) as the HT generator and BM25 and NVIDIA NV-Embed-2 Lee et al. (2025) as the retrievers. Each model is evaluated with inference mode using both TND/QTND settings. A different dense retriever is used here to gauge the universality of our approach. We use the code split of ToolRet for these tests. The results are presented in Table 3. As we see, ToolDreamer is flexible enough to work with different LLMs with hardly any loss of performance. Thus, there is no necessity to stick to API models for HT generation, which shows the cost-benefit of using ToolDreamer.
@@
4.3.2 Weak Tool Alignment
Alignment Method NDCG@10 P@10 R@10 MRR
Original 39.79 9.16 53.46 40.04
DPR/Hungarian 39.48-1% 9.10-.7% 53.0-.9% 39.81-.6%
Qwen/Greedy 39.7-.2% 9.15-.1% 53.22-.4% 40.0-.1%
4.3.3 LLM List Fusion
Table 5 shows the results of this test. As seen, LLM reranking provides additional improvements over RRF. This makes sense as RRF is a simple statistical technique which simply considers the overall ranking of elements across lists, whereas an LLM does deeper reasoning to figure out which tools actually make sense for a query. However, these results come with caveats such as additional API costs, ignoring instructions, and hallucinations (spurious generations) Huang et al. (2025), which we elaborate in App. D. Considering these issues, we prioritise RRF over LLM-ranking in ToolDreamer as it offers reproducible and deterministic guarantees. That said, it is easy to equip ToolDreamer with LLM-ranking, should cost/instruction-refusal not be a pressing concern.
Approach NDCG@10 P@10 R@10 MRR
TND/RRF 39.79 9.16 53.46 40.04
QTND/RRF 42.47 9.34 54.26 43.88
TND/LLM-F 46.48 10.16+9% 57.64 46.83
QTND/LLM-F 46.53+10% 10.14 57.67+6% 46.89+7%
Table 5: Impact of using different ranked-list fusion methods. RRF = Reciprocal Rank Fusion; LLM-F = Large Language Model (GPT-4.1) Fusion. All scores are from the trained Qwen3 model using TND/QTND vectors. Original scores are given for easier comparison. Improvements are shown against best baseline (QTND).
@@ 5 Conclusion
We identify two limitations with our framework. First, our framework requires high-quality hypothetical tools to work well (§4.3.1). This requires a few rounds of testing to standardise the prompt, by experimenting with format, instructions and examples. However, considering the importance of prompt quality Long et al. (2025) for most generative tasks, we find this requirement reasonable for ToolDreamer. Second, compared to regular predictions, reciprocal rank fusion adds another layer of processing to increase the evaluation time complexity. However, from our evaluations, we did not find this to be a significant bottleneck considering the capability of modern hardware.
@@ Appendix A Benchmarking Current Retrievers
  • Semi-Sparse: These are neural models that produce high-dimension vectors (mostly 0’s) but learn the non-zero elements to capture better semantics while optimising for search efficiency. We select SPLADE Lassance et al. (2024) and opensearch Geng et al. (2024) for this category
  • Dense: Models that generate low-dimension embeddings aimed at capturing deeper semantics over simple token matching. These include DPR, BGE Xiao et al. (2024), Jina Günther et al. (2023), MiniLM Wang et al. (2021), NV-Embed, LGAI-Embedding Choi et al. (2025), QZhou-Embedding Yu et al. (2025) and Qwen3 Zhang et al. (2025b). We only use one proprietary model, i.e., OpenAI’s text-embedding-3-small OpenAI (2024) to maintain cost.
@@ Appendix B LLM Performance v/s Tool Number
To highlight the issue of tool limits, we examine how GPT-4.1 OpenAI (2025) and Qwen3’s Yang et al. (2025) performance changes as the number of tools increases (c.f. Fig. 3). We use SealTools Wu et al. (2025), a dataset for LLM-tool call training for illustration. For each query, we randomly sample negative (irrelevant) tools to meet the tool counts (10-128). For example, if a query has 1 gold tool, we randomly sample 9 negative tools and shuffle their order to prevent location bias Liu et al. (2024). As we can see, at relatively small scales (10-128 tools), each model handles tool calling relatively well, showing consistent performance across varying tool counts. However, beyond 50 (for Qwen3) and 128 (for GPT-4.1), each model hits their limit (Qwen: 32K context window; GPT: 1M), being unable to register more tools. GPT has another issue - since it is an API-based model, OpenAI/Azure restricts the maximum number of tools to 128 Microsoft (2025). Thus, despite theoretically being able to support even more tools, given a 1M context window, GPT cannot be provided with more tools. These issues reinforce the necessity for tool retrieval.
@@ Appendix E Dataset Filtration
The provided training dataset has ~200K samples. However, many samples had profanities, repeated queries, negative tools appearing in the positive tools list and questions with 0 negative tools (after removing the overlapping tools). To keep our data generation reasonable and devoid of such entries, we filter and sample 5K unique queries from the base set.

### v2

@@ 4.1 Experiment Settings
Evaluation Metrics. We evaluate retriever quality using 4 standard IR (Information Retrieval) metrics. These are (i) NDCG@K/N@K (Normalized Discounted Cumulative Gain): evaluates the quality of a ranked list by rewarding relevant items being placed at the top of the list and penalizing them being placed towards the bottom; (ii) P@K (Precision): The fraction of retrieved items that are relevant to the search; (iii) R@K (Recall): The fraction of relevant items that were retrieved, and (iv) MRR (Mean Reciprocal Rank): Inverse rank of the first relevant item in the retrieved list. For each measure, @K means the number of items that are retrieved. Following ToolRet, we set K to 10.
Baselines. In §2, we mention various related approaches for tool-retriever training. However, Xu et al. (2024) does not have publicly available code. As such, we use (i) Shi et al. (2025) (ToolRet (TR)), who train their models with both query-tool and (query + instruction) - tool similarity. However, we intentionally avoid incorporating the instructions to ensure a consistent evaluation with existing benchmarks which lack such annotations as our goal is to evaluate ToolDreamer under the minimal assumptions that broadly apply across tool datasets; (ii) COLT Qu et al. (2024), a complex two-step framework that first tunes a model (Contriever Izacard et al. (2022)) using standard InfoNCE with query-tool similarity, followed by further optimisation with a graph neural network (GNN) using various decompositions of the training dataset; (iii) use HT directly for retrieval similar to Kachuee et al. (2025) (K in Table 1).
HT Generation Model. We use GPT-4.1 OpenAI (2025) to generate hypothetical tools for both training and test data queries. Additionally, we show that ToolDreamer can also use open-source models for HT generation (§4.4). Thus, it is not restricted by the choice of LLM used.
4.2 ToolDreamer Performance Analysis (RQ1)
To evaluate the tool retrieval performance of ToolDreamer, we choose a dense Qwen3-8B and a sparse BM25 model (App. A). Table 1 shows the results of ToolDreamer against related baselines.
@@ 4.2 ToolDreamer Performance Analysis (RQ1)
Next, we see how strong our framework is against COLT, another related baseline. Firstly, the results from phase-1 training (general query-tool alignment) are far superior to phase-2 (GNN-optimisation). This makes sense as they train their model using the standard query-tool InfoNCE objective, which works relatively well. However, the increased complexity of grouping tools, aligning them with queries, and using a GNN for further optimisation plummets performance. In this regard, ToolDreamer is a much simpler and scalable approach capable of handling a variety of questions.
Finally, we would like to clarify that in this work, we intentionally exclude the instruction field of ToolRet dataset with the aim of developing/evaluating ToolDreamer without assuming any a priori instruction directives. Note that most existing tool-calling datasets lack instruction field annotations and annotating query instructions for new real-world use cases can be resource-intensive (e.g., requiring human experts to manually craft seed instructions as in ToolRet). Without dependency on pre-exisitng instructions, ToolDreamer can be readliy applied to other datasets as well as to new tool-calling use cases.
4.3 Ablation Studies (RQ2)
We conduct various ablation studies to investigate the importance of each step of ToolDreamer: (i) the impact of HT quality (§4.3.1) (ii) using a weak discriminator for tool alignment along with the alignment algorithm itself (§4.3.2), and (iii) whether using an LLM for ranked list fusion offers additional benefits or not (§4.3.3).
@@ 4.3.1 Hypothetical Tool Quality
4.3.2 Weak Tool Alignment
Alignment Method N@10 P@10 R@10 MRR
Original 39.79 9.16 53.46 40.04
DPR/Hungarian 39.48-1% 9.10-.7% 53.0-.9% 39.81-.6%
Qwen/Greedy 39.70-.2% 9.15-.1% 53.22-.4% 40.0-.1%
4.3.3 LLM List Fusion
Table 4 shows the results of this test. As seen, LLM reranking provides additional improvements over RRF. This makes sense as RRF is a simple statistical technique that simply considers the overall ranking of elements across lists, whereas an LLM does deeper reasoning to figure out which tools actually make sense for a query. However, these results come with caveats such as additional API costs, ignoring instructions, and hallucinations (spurious generations) Huang et al. (2025), which we elaborate in App. D. Considering these issues, we prioritise RRF over LLM-ranking in ToolDreamer as it offers reproducible and deterministic guarantees. That said, it is easy to equip ToolDreamer with LLM-ranking, should cost/instruction-refusal not be a pressing concern.
Approach NDCG@10 P@10 R@10 MRR
TND/RRF 39.79 9.16 53.46 40.04
QTND/RRF 42.47 9.34 54.26 43.88
TND/LLM-F 46.48 10.16+9% 57.64 46.83
QTND/LLM-F 46.53+10% 10.14 57.67+6% 46.89+7%
Table 4: Impact of using different ranked-list fusion methods. LLM-F = LLM (GPT-4.1) Fusion. All scores are from the trained Qwen3 model using TND/QTND vectors. Improvements are shown against the best baseline (QTND).
4.4 Varying HT Generator (RQ3)
In our experiments, we use GPT-4.1 to create our HT’s. However, this might be a bottleneck in cases where API-calling might be an issue, either due to cost or privacy concerns Yao et al. (2024). Thus, we wanted to see how switching out GPT-4.1 for an open-source model impacted performance. To this end, we use Qwen3-32B (with the same prompt (Fig. 8) as GPT-4.1) as the HT generator and BM25 and NVIDIA NV-Embed-2 Lee et al. (2025) as the retrievers. Each model is evaluated with inference mode (no training) using both TND/QTND settings. A different dense retriever is used here to gauge the universality of our approach. We use the code split of ToolRet for these tests. The results are presented in Table 5. As we see, ToolDreamer is flexible enough to work with different LLMs without much loss of performance. Thus, there is no necessity to stick to API models for HT generation, which shows the cost-benefit of using ToolDreamer.
@@ 4.4 Varying HT Generator (RQ3)
4.5 Production Considerations (RQ4)
As we report our main scores using an API-based model, it is necessary to understand the associated costs, as higher monetary requirements can make the framework unsuitable for real-time production demands. Implementing ToolDreamer cost us less than \$5 using OpenAI’s batch mode and the cheaper GPT-4.1. Thus, we can claim that our framework is highly cost-effective.
To gauge wall-time, we consider the query Can you tell me about the historical events of April 21st? from the dataset. Basic retrieval (Qwen) takes ~0.04 seconds. With ToolDreamer, it takes us ~8 seconds to generate the hypothetical tools using Qwen3-32B (on 1 A100 card) and ~2.5 seconds with GPT. Semantic search + RRF takes an additional 0.5 seconds, bringing the total to ~3 - 8.5 seconds. While this does introduce some latency, it is unlikely that in production, clients will constantly be asking questions that require tool usage, which helps justify the time constraints.
@@ 5 Conclusion
We identify two limitations with our framework and offer potential workarounds. First, ToolDreamer requires high-quality hypothetical tools to work well (§4.3.1). This requires a few rounds of testing to standardise the prompt, by experimenting with format, instructions, and examples. However, considering the importance of prompt quality Long et al. (2025) for most generative tasks, we find this requirement reasonable for ToolDreamer. Second, compared to regular retrieval, RRF adds another layer of processing to increase the evaluation time complexity. However, from our evaluations (§4.5), we did not find this to be a significant bottleneck considering the capability of modern hardware.
@@ Appendix A Benchmarking Current Retrievers
  • Semi-Sparse: These are neural models that produce high-dimensional vectors (mostly 0’s) but learn the non-zero elements to capture better semantics while optimising for search efficiency. We select SPLADE Lassance et al. (2024) and opensearch Geng et al. (2024) for this category
  • Dense: Models that generate low-dimensional embeddings aimed at capturing deeper semantics over simple token matching. These include DPR, BGE Xiao et al. (2024), Jina Günther et al. (2023), MiniLM Wang et al. (2021), NV-Embed, LGAI-Embedding Choi et al. (2025), QZhou-Embedding Yu et al. (2025) and Qwen3 Zhang et al. (2025b). We only use one proprietary model, i.e., OpenAI’s text-embedding-3-small OpenAI (2024) to maintain cost.
@@ Appendix B LLM Performance v/s Tool Number
To highlight the issue of tool limits, we examine how GPT-4.1 OpenAI (2025), and Qwen3’s Yang et al. (2025) performance changes as the number of tools increases (c.f. Fig. 3). We use SealTools Wu et al. (2025), a dataset for LLM-tool call training for illustration. For each query, we randomly sample negative (irrelevant) tools to meet the tool counts (10-128). For example, if a query has 1 gold tool, we randomly sample 9 negative tools and shuffle their order to prevent location bias Liu et al. (2024). As we can see, at relatively small scales (10-128 tools), each model handles tool calling relatively well, showing consistent performance across varying tool counts. However, beyond 50 (for Qwen3) and 128 (for GPT-4.1), each model hits their limit (Qwen: 32K context window; GPT: 1M), being unable to register more tools. GPT has another issue - since it is an API-based model, OpenAI/Azure restricts the maximum number of tools to 128 Microsoft (2025). Thus, despite theoretically being able to support even more tools, given a 1M context window, GPT cannot be provided with more tools. These issues reinforce the necessity for tool retrieval.
@@ Appendix E Dataset Filtration
The provided training dataset has ~200K samples. However, many samples had profanities, repeated queries, negative tools appearing in the positive tools list, and questions with 0 negative tools (after removing the overlapping tools). To keep our data generation reasonable and devoid of such entries, we filter and sample 5K unique queries from the base set.
@@ 2. Tool Design
Appendix H Paper Code
The code for this paper is publicly available at https://github.com/boschresearch/ToolDreamer.